OpenStudy (anonymous):

Helpppp.A density curve for all the possible ages between 0 years and 100 years is in the shape of a triangle. What is the height of the triangle? A. 0.08 B. 0.04 C. 0.01 D. 0.02

OpenStudy (amistre64):

since the area under any density curve is 1 do you recall the formula for the area of a triangle?

OpenStudy (anonymous):

Not so much,i have the hardest time with these

OpenStudy (amistre64):

a triangle is half a rectangle, and the area of a rectangle is the product of its base and height

OpenStudy (amistre64):

A = bh/2 A = 1, b = 100 solve for h

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

im not sure what to start plugging in

OpenStudy (amistre64):

plug in what is known .... then see whats left to solve

OpenStudy (amistre64):

or, shuffle for h, then plug in

OpenStudy (amistre64):

A = bh/2 2A = bh 2A/b = h

OpenStudy (anonymous):

i keep getting whole numbers

OpenStudy (amistre64):

2(1)/100 = h = 2/100

OpenStudy (anonymous):

ok i got 0.02

OpenStudy (amistre64):

thatll do it

OpenStudy (anonymous):

thanks so much friend !

OpenStudy (amistre64):

youre welcome